OpenStudy (anonymous):

need help with finding an answer to: 5t-1=-11 the text sais to solve the equation

OpenStudy (anonymous):

you would do \[5t = -11 + 1\]\[5t = -10\]\[t = \frac{-10}{5} = -2\]

OpenStudy (anonymous):

-2 = t

OpenStudy (anonymous):

thank you